If you react 5.625 L of H2S in excess oxygen, how many grams of water will be produced


1
Expert's answer
2022-04-08T14:47:04-0400

Solution:

At STP, one mole of any gas occupies a volume of 22.4 L.

Therefore,

Moles of H2S = (5.625 L H2S) × (1 mol H2S / 22.4 L H2S) = 0.2511 mol H2S


Balanced chemical equation:

2H2S + 3O2 → 2SO2 + 2H2O

According to stoichiometry:

2 mol of H2S produce 2 mol of H2O

Thus, 0.2511 mol of H2S produce:

(0.2511 mol H2S) × (2 mol H2O / 2 mol H2S) = 0.2511 mol H2O


The molar mass of water (H2O) is 18.0153 g/mol

Therefore,

Mass of H2O = (0.2511 mol H2O) × (18.0153 g H2O / 1 mol H2O) = 4.524 g H2O

Mass of H2O = 4.524 g


Answer: 4.524 grams of water (H2O) will be produced